# On the Classification of Dillon’s APN Hexanomials

### v1

@@ Abstract.
In this paper, we undertake a systematic analysis of a class of hexanomial functions over finite fields of characteristic 2 proposed by Dillon in 2006 as potential candidates for almost perfect nonlinear (APN) functions, pushing the analysis a lot further than what has been done via the partial APN concept in (Budaghyan et al., DCC 2020). These functions, defined over $\mathbb{F}_{q^{2}}$ where $q=2^{n}$, have the form
@@
Using algebraic number theory and methods on algebraic varieties over finite fields, we establish necessary conditions on the coefficients $A,B,C,D,E$ that must hold for the corresponding function to be APN. Our main contribution is a comprehensive case-by-case analysis that systematically excludes large classes of Dillon’s hexanomials from being APN based on the vanishing patterns of certain key polynomials in the coefficients. Through a combination of number theory, algebraic-geometric techniques and computational verification, we identify specific algebraic obstructions—including the existence of absolutely irreducible components in associated varieties and degree incompatibilities in polynomial factorizations—that prevent these functions from achieving optimal differential uniformity. Our results significantly narrow the search space for new APN functions within this family and provide a theoretical roadmap applicable to other classes of potential APN functions. We complement our theoretical work with extensive computations. Through exhaustive searches on $\mathbb{F}_{2^{2}}$ and $\mathbb{F}_{2^{4}}$ and random sampling on $\mathbb{F}_{2^{6}}$ and $\mathbb{F}_{2^{8}}$, we identified thousands of APN hexanomials. Subsequent classification based on CCZ-invariants reveals a large number of inequivalent classes, many of which are not CCZ-equivalent to the known Budaghyan-Carlet family (Budaghyan-Carlet, IEEE Trans. Inf. Th., 2008).
@@
Finite fields, permutation polynomials, APN, varieties, irreducible components
@@ 1. Introduction and motivation
The security of a block cipher depends upon the immunity of its substitution boxes against many cryptographic attacks. For example, a low differential uniformity [13] is needed in order to resist the differential attacks [3]. For a positive integer $n>0$, the differential uniformity of an $(n,n)$-function $F:\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ is defined as the maximum number of solutions $x\in\mathbb{F}_{p^{n}}$ of the differential equation $F(x+a)+F(x)=b$, where $a\neq 0,b\in\mathbb{F}_{2^{n}}$. The lowest possible differential uniformity of functions over finite fields of even characteristic is $2$ and such functions are called almost perfect nonlinear (APN).
@@
Our approach transforms this problem into the study of algebraic varieties over finite fields. By reformulating the APN condition as a question about the existence of certain algebraic varieties and their irreducible components, we can apply powerful tools from algebraic geometry to obtain results that would be difficult or impossible to achieve through direct computational methods alone. This geometric perspective not only provides theoretical insights but also leads to practical algorithms for determining when specific instances of Dillon’s hexanomials fail to be APN.
@@ 2. A key theorem
Suppose that there exists a variety $\mathcal{C}$ contained in $\mathcal{W}$ that is absolutely irreducible and fixed by $\phi$, where $\phi(A,B,C,D,E,X_{0},X_{1},Z_{0},Z_{1})=(A^{q},B^{q},C^{q},D^{q},E^{q},X_{1},X_{0},Z_{1},Z_{0})$ and not contained in the hyperplanes $X_{0}=0$, $X_{1}=0$, $Z_{0}=X_{0}$, $Z_{1}=X_{1}$, $Z_{0}=0$, $Z_{1}=0$. Then, if $q$ is large enough $f_{A,B,C,D,E}(x)$ is not APN. In particular, if the dimension of $\mathcal{C}$ is $2$ and $q\geq 2^{20}$, then $f_{A,B,C,D,E}(x)$ is not APN. Conversely, if $\mathcal{W}$ is contained in the union of the forbidden hyperplanes $\pi_{i}$, $i=1,\ldots,6$, defined by $X_{0}=0$, $X_{1}=0$, $Z_{0}=X_{0}$, $Z_{1}=X_{1}$, $Z_{0}=0$, $Z_{1}=0$, then $f_{A,B,C,D,E}(x)$ is APN.
@@ 2. $(C2)$
    Since $H(X_{0},X_{1},Z_{0},1)=X_{1}^{2}+X_{1}+AX_{0}Z_{0}(X_{0}+Z_{0})$ has constant term (in $X_{1}$) of degree three in $X_{0}$ e $Z_{0}$, $H(X_{0},X_{1},Z_{0},1)$ and thus $H(X_{0},X_{1},Z_{0},Z_{1})$ is absolutely irreducible.
@@ 3. Case $B=0$
    If $(A^{q+1}+1)=0$, $(C^{q+1}+1)\neq 0$, $AE^{q}+E\neq 0$, and $T^{3}+CT^{2}+AC^{q}T+A$ has a root $k\in\mathbb{F}_{q^{2}}$ with $k^{q+1}=1$, then $f_{A,B,C,D,E}(x)$ is not APN.
@@
The APN status of the function remains open in the following cases:
  • When $(A^{q+1}+1)=0$, $(C^{q+1}+1)\neq 0$, $AE^{q}+E\neq 0$, and $T^{3}+CT^{2}+AC^{q}T+A$ has no roots in $\mathbb{F}_{q^{2}}$, the function may be APN if the conic $AE^{q}Z_{0}^{2}+CZ_{0}+Z_{1}=0$ contributes only trivial solutions. This requires further analysis of the $\mathbb{F}_{q}$-rational points on this conic.
  • When $(A^{q+1}+1)\neq 0$, $(C^{q+1}+1)=0$, and $AE^{q}+A^{q}C^{3}E+C^{3}E^{q}+E=0$, the polynomial $H$ becomes absolutely irreducible. The APN status of such functions remains undetermined and requires detailed geometric analysis beyond the scope of this paper.
@@
Proof of Part (1): Assume $(A^{q+1}+1)(C^{q+1}+1)\neq 0$. First, we verify that $AE^{q}+E\neq 0$ under our hypotheses. Suppose, for contradiction, that $AE^{q}+E=0$. If $E=0$, this is satisfied trivially. If $E\neq 0$, then $AE^{q}=E$ implies $A^{q+1}=1$. Combined with our hypotheses $B=0$, $D=AC^{q}$ (from $AC^{q}+D=0$), and $E^{q}=A^{q}E$ (from $AE^{q}=E$), we have $A^{q+1}=1$, $D=AC^{q}$, $B^{q}=A^{q}B$ (trivially, since $B=0$), and $E^{q}=A^{q}E$. If additionally $C\neq 0$ and $D\neq 0$, these are precisely the conditions for $(C2)$ from Proposition 2.3, contradicting our hypothesis that $(C2)$ does not hold.
If $C=0$, then $D=0$ (from $D=AC^{q}$), and we have $A\neq 0$ (since $A^{q+1}=1$), $C=D=0$, $B^{q}=A^{q}B$, and $E^{q}=A^{q}E$, which are precisely the conditions for $(C1)$, contradicting our hypothesis that $(C1)$ does not hold. Therefore, we must have $AE^{q}+E\neq 0$. Now, by Remark 2.6, if $b_{2}X_{0}^{2}+b_{1}X_{0}+b_{0}$ splits into degree-one factors in $X_{0}$, then
@@
Proof of Part (2): Assume $(A^{q+1}+1)=0$, $(C^{q+1}+1)\neq 0$, $AE^{q}+E\neq 0$, and the cubic $T^{3}+CT^{2}+AC^{q}T+A$ has a root $k\in\mathbb{F}_{q^{2}}$ with $k^{q+1}=1$. In this case, Equation (2.3) becomes
@@
By hypothesis, this cubic in $T$ has a root $k\in\mathbb{F}_{q^{2}}$ with $k^{q+1}=1$. Since $k^{q+1}=1$, we have $k^{q}=k^{-1}$, which means the line $\mathcal{L}_{k}$ defined by $Z_{1}=kZ_{0}$ is $\mathbb{F}_{q}$-rational.
@@ 4. Case $B\neq 0$
Suppose that conditions $(C1)$ and $(C2)$ do not hold. Suppose that $h_{1}=0$ and $BC^{q}+B^{q}D\neq 0.$ Let $q$ be large enough. If $f_{A,B,C,D,E}(x)$ is APN then one of the following conditions possibly holds
@@ 5. $B\neq 0$ and (C6) does not hold
Solutions of $\overline{G}(X_{0},Z_{0},Z_{1})=0$ not contained in the forbidden hyperplanes can arise only from the factors $AD^{q}Z_{0}+AE^{q}Z_{0}^{2}+Z_{1}$ and $Z_{1}^{3}+AD^{q}Z_{0}Z_{1}^{2}+DZ_{0}^{2}Z_{1}+AZ_{0}^{3}$. Let us consider the factor
@@ 6. Computational verification and discovery of new APN classes
To assess the diversity of discovered functions, we classified them using CCZ-invariants: differential uniformity, differential spectrum, Walsh spectrum distribution, and nonlinearity. Functions sharing identical invariant tuples were grouped into classes. Since distinct CCZ-equivalence classes may share the same invariants, our class counts provide lower bounds on the true number of inequivalent classes. For $q=16$, where we found relatively few APN functions, we performed complete pairwise CCZ-equivalence testing to obtain exact counts.
@@
    For $q=2$ (exhaustive): $390$ APN functions in at least $10$ distinct CCZ-invariant classes. Of these, $2$ are CCZ-equivalent to the Budaghyan-Carlet (BC) family [5], leaving at least $8$ new classes outside of the BC family.
@@
    For $q=4$ (exhaustive): $28,170$ APN functions in at least $182$ distinct classes, with $1$ BC-equivalent, leaving at least $181$ new classes outside of the BC family.
@@
    For $q=8$ ($60,000$ random candidates): $104$ APN functions in at least $101$ distinct CCZ-equivalent classes, all outside of the BC family.
@@
    For $q=16$ ($120,000$ random candidates): $25$ APN functions in exactly $2$ distinct CCZ-equivalence classes (verified by complete pairwise testing).
Representatives for each class appear in Tables 617. Remarkably, none of the discovered APN functions are permutations.
@@
The enumeration and classification were performed using SageMath implementations of the algorithms described above. Complete computational code and output files are available at [14]. ∎
@@
Yet within this rarefied landscape, we find remarkable diversity. The discovery of at least $60$ new CCZ-invariant classes for $q=4$ demonstrates that Dillon’s hexanomial family is significantly richer than previously recognized. While the Budaghyan-Carlet construction [5] established that this family contains APN functions, our results show it contains many inequivalent classes. For $q=2$, 80% of classes are not BC-equivalent; for $q=4$, this rises to 93.75%; for $q=8$, it reaches 91.67%. This validates Dillon’s 2006 intuition that hexanomials merit systematic investigation.
The universal absence of permutations among all tested APN hexanomials is particularly striking. This distinguishes Dillon’s family from other APN constructions where permutation polynomials exist, suggesting these hexanomials possess structural features fundamentally incompatible with bijectivity. Understanding this phenomenon could provide insight into the relationship between the APN property and injectivity in polynomial mappings over finite fields.

### v3

@@ Abstract.
In this paper, we undertake a systematic analysis of a class of hexanomial functions over finite fields of characteristic 2 proposed by Dillon in 2006 as potential candidates for almost perfect nonlinear (APN) functions, pushing the analysis a lot further than what has been done via the partial APN concept in Budaghyan, Kaleyski, Riera, and Stănică (Des. Codes Cryptogr. 88, 2020), and Chase and Lisoněk (Cryptogr. Commun. 13, 2021). These functions, defined over $\mathbb{F}_{q^{2}}$ where $q=2^{n}$, have the form
@@
Using algebraic number theory and methods on algebraic varieties over finite fields, we establish necessary conditions on the coefficients $A,B,C,D,E$ that must hold for the corresponding function to be APN. Our main contribution is a comprehensive case-by-case analysis that systematically excludes large classes of Dillon’s hexanomials from being APN based on the vanishing patterns of certain key polynomials in the coefficients. Through a combination of number theory, algebraic-geometric techniques and computational verification, we identify specific algebraic obstructions—including the existence of absolutely irreducible components in associated varieties and degree incompatibilities in polynomial factorizations—that prevent these functions from achieving optimal differential uniformity. Our results significantly narrow the search space for new APN functions within this family and provide a theoretical roadmap applicable to other classes of potential APN functions. We complement our theoretical work with extensive computations. Through exhaustive searches on $\mathbb{F}_{2^{2}}$ and $\mathbb{F}_{2^{4}}$ and random sampling on $\mathbb{F}_{2^{6}}$ and $\mathbb{F}_{2^{8}}$, we identified hundreds of APN hexanomials. Our classification, based on complete CCZ-equivalence testing, reveals that while many instances exist, they belong to very few distinct classes. For $q\in\{2,4\}$, all discovered APN functions are CCZ-equivalent to the known Budaghyan-Carlet (BC) family (Budaghyan-Carlet, IEEE Trans. Inf. Th., 2008). However, for larger dimensions, none of them seem to be equivalent to the BC family.
@@
Finite fields, permutation polynomials, varieties, irreducible components
@@ 1. Introduction and motivation
The security of a block cipher depends upon the immunity of its substitution boxes against many cryptographic attacks. For example, a low differential uniformity [10] is needed in order to resist the differential attacks [1]. For a positive integer $n>0$, the differential uniformity of an $(n,n)$-function $F:\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ is defined as the maximum number of solutions $x\in\mathbb{F}_{2^{n}}$ of the differential equation $F(x+a)+F(x)=b$, where $a\neq 0,b\in\mathbb{F}_{2^{n}}$. The lowest possible differential uniformity of functions over finite fields of even characteristic is $2$ and such functions are called almost perfect nonlinear (APN).
@@
Our approach transforms this problem into the study of algebraic varieties over finite fields. Janwa and Wilson were the first to employ algebraic varieties to study APN functions, see [8]. For a deeper introduction to algebraic varieties we refer the interested reader to [7]. By reformulating the APN condition as a question about the existence of certain algebraic varieties and their irreducible components, we can apply powerful tools from algebraic geometry to obtain results that would be difficult or impossible to achieve through direct computational methods alone. This geometric perspective not only provides theoretical insights but also leads to practical algorithms for determining when specific instances of Dillon’s hexanomials fail to be APN.
@@ 2. A key theorem
Recall that the APN condition requires that for $a\neq 0$, the only solutions to the derivative equation are $x=0$ and $x=a$. In terms of our coordinates $X_{0},X_{1},Z_{0},Z_{1}$, these trivial solutions correspond to specific hyperplanes in the affine space $\mathbb{A}^{4}(\mathbb{F}_{q^{2}})$. Additionally, degenerate cases where $a$ is not a valid parameter correspond to $Z_{0}=0$ or $Z_{1}=0$. We formalize this locus as follows.
Definition 2.1 (Forbidden planes).
The forbidden locus, denoted by $\Pi$, is the union of the six hyperplanes $\pi_{1},\dots,\pi_{6}$ in $\mathbb{A}^{4}(\mathbb{F}_{q^{2}})$ defined by the trivial APN solutions and the degenerate parameters:
$\pi_{1}$ $:X_{0}=0,$ $\pi_{2}$ $:X_{1}=0,$
$\pi_{3}$ $:X_{0}+Z_{0}=0,$ $\pi_{4}$ $:X_{1}+Z_{1}=0,$
$\pi_{5}$ $:Z_{0}=0,$ $\pi_{6}$ $:Z_{1}=0.$
Consequently, $f_{A,B,C,D,E}$ is APN if and only if the set of $\mathbb{F}_{q}$-rational points of the variety $\mathcal{W}$ is contained in $\Pi$.
@@
Suppose that there exists a variety $\mathcal{C}$ contained in $\mathcal{W}$ that is absolutely irreducible and fixed by $\phi$, where $\phi(A,B,C,D,E,X_{0},X_{1},Z_{0},Z_{1})=(A^{q},B^{q},C^{q},D^{q},E^{q},X_{1},X_{0},Z_{1},Z_{0})$ and not contained in the hyperplanes $X_{0}=0$, $X_{1}=0$, $Z_{0}=X_{0}$, $Z_{1}=X_{1}$, $Z_{0}=0$, $Z_{1}=0$. Then, if $q$ is large enough $f_{A,B,C,D,E}(x)$ is not APN. In particular, if the dimension of $\mathcal{C}$ is $2$ and $q\geq 2^{20}$, then $f_{A,B,C,D,E}(x)$ is not APN. Conversely, if $\mathcal{W}$ is contained in $\Pi$, then $f_{A,B,C,D,E}(x)$ is APN.
@@
Remark 2.4.
The lower bound $q\geq 2^{20}$ established in Theorem 2.3 is a sufficient condition derived from the worst-case error terms in the Lang-Weil bound (using for instance the Cafure-Matera extimates). This ensures that the number of rational points on the variety $\mathcal{C}$ strictly exceeds the number of points contained in the forbidden hyperplanes. However, from a geometric perspective, the existence of an absolutely irreducible component $\mathcal{C}\not\subseteq\bigcup\pi_{i}$ is a structural obstruction to the APN property that typically manifests for much smaller $q$. In the vast majority of cases, such a variety will possess $\mathbb{F}_{q}$-rational points outside the forbidden locus long before the asymptotic lower bound is reached. This is strongly supported by our computational classification in Section 6, where we observe that for $q\in\{2,4,8,16\}$, the function $f_{A,B,C,D,E}$ fails to be APN when the geometric conditions for the existence of such a component are met.
@@ 2. $(C2)$
    Since $H(X_{0},X_{1},Z_{0},1)=X_{1}^{2}+X_{1}+AX_{0}Z_{0}(X_{0}+Z_{0})$ has constant term (in $X_{1}$) of degree three in $X_{0}$ and $Z_{0}$, $H(X_{0},X_{1},Z_{0},1)$ and thus $H(X_{0},X_{1},Z_{0},Z_{1})$ is absolutely irreducible.
@@ 3. Case $B=0$
    If $(A^{q+1}+1)=0$, $(C^{q+1}+1)\neq 0$, $AE^{q}+E\neq 0$, and $T^{3}+CT^{2}+AC^{q}T+A$ has a root $k\in\mathbb{F}_{q^{2}}$ with $k\neq 1$ and $k^{q+1}=1$, then $f_{A,B,C,D,E}(x)$ is not APN.
@@
The APN status of the function remains open when $(A^{q+1}+1)=0$, $(C^{q+1}+1)\neq 0$, $AE^{q}+E\neq 0$, and $T^{3}+CT^{2}+AC^{q}T+A$ has no roots in $\mathbb{F}_{q^{2}}$, the function may be APN if the conic $AE^{q}Z_{0}^{2}+CZ_{0}+Z_{1}=0$ contributes only trivial solutions. This requires further analysis of the $\mathbb{F}_{q}$-rational points on this conic.
@@
Proof of Part (1): Assume $(A^{q+1}+1)(C^{q+1}+1)\neq 0$.
First, assume $AE^{q}+E=0$. Note that $C\neq 0$ otherwise we are in case $(C1)$ or in the trivial case $A=B=C=D=E=0$. Now, by Remark 2.8, if $b_{2}X_{0}^{2}+b_{1}X_{0}+b_{0}$ splits into degree-one factors in $X_{0}$, then
with $\beta^{2}=b_{2}$ and $\gamma(\gamma+\beta Z_{0})=b_{0}$. Since $\deg(\beta)=2$ and $\deg(b_{0})=6$, then $\gamma$ is homogeneous of degree 3. Put
$\gamma=rZ_{0}^{3}+sZ_{0}^{2}Z_{1}+tZ_{0}Z_{1}^{2}+uZ_{1}^{3}.$
Such an $\gamma$ must make $h:=\gamma^{2}+\gamma\beta Z_{0}+b_{0}$ the zero polynomial in $Z_{0}$ and $Z_{1}$. Let $h:=\sum_{i=0}^{6}h_{i}Z_{1}^{i}Z_{0}^{6-i}$, where
$h_{0}$ $=$ $r\Big(r+(A^{q+1}+1)C\Big),$
$h_{1}$ $=$ $(A^{q+1}+1)\Big(r(C^{q+1}+1)+sC\Big),$
$h_{2}$ $=$ $rC^{q}(A^{q+1}+1)+s^{2}+s(A^{q+1}+1)(C^{q+1}+1)+tC(A^{q+1}+1)+A^{q+1}C^{q+1}(C^{q+1}+1),$
$h_{3}$ $=$ $(A^{q+1}+1)\Big(sC^{q}+t(C^{q+1}+1)+uC+C^{q}(C^{q+1}+1)\Big),$
$h_{4}$ $=$ $t^{2}+tC^{q}(A^{q+1}+1)+u(A^{q+1}+1)(C^{q+1}+1),$
$h_{5}$ $=$ $C^{q}(A^{q+1}+1)u,$
$h_{6}$ $=$ $u^{2}.$
From $h_{1}=h_{3}=h_{5}=h_{6}=0$, we get $r=sC(C^{q+1}+1)^{-1}$, $t=C^{q}(C^{q+1}+1)^{-1}(s+C^{q+1}+1)$, $u=0$. After substituting them in $h_{4}$, from $h_{4}=0$ we get either $s=C^{q+1}+1$ or $s=A^{q+1}(C^{q+1}+1)$. In both cases, a contradiction follows from $h_{0}=0$.
Now, assume $AE^{q}+E\neq 0$. By Remark 2.8, if $b_{2}X_{0}^{2}+b_{1}X_{0}+b_{0}$ splits into degree-one factors in $X_{0}$, then
@@
Proof of Part (2): Assume $(A^{q+1}+1)=0$, $(C^{q+1}+1)\neq 0$, $AE^{q}+E\neq 0$, and the cubic $T^{3}+CT^{2}+AC^{q}T+A$ has a root $k\in\mathbb{F}_{q^{2}}$ with $k\neq 1$ and $k^{q+1}=1$. In this case, Equation (2.3) becomes
@@
By hypothesis, this cubic in $T$ has a root $k\in\mathbb{F}_{q^{2}}$ with $k\neq 1$ and $k^{q+1}=1$. Since $k^{q+1}=1$, we have $k^{q}=k^{-1}$, which means the line $\mathcal{L}_{k}$ defined by $Z_{1}=kZ_{0}$ is $\mathbb{F}_{q}$-rational.
@@
The case $B=E=0$ was previously investigated [9], where the authors characterize APN functions with $A,D\in\mathbb{F}_{q^{2}}$ and $C\in\mathbb{F}_{q}$, see Theorem 1.2. Subsequently, Chase and Lisoněk in [5] show that if $A,D,C\in\mathbb{F}_{q^{2}}$, any such APN functions are necessarily equivalent to Gold functions. The characterization result mentioned above is not fully correct; indeed, we found a counterexample. If we consider $q^{2}=2^{8}$ and $\alpha$ a root of the polynomial $x^{8}+x^{4}+x^{3}+x^{2}+1$, then $F(x)=x^{3q}+x^{2q+1}+\alpha^{8}x^{q+2}+\alpha x^{3}$ is not APN since the equation $F(x+1)+F(x)=0$ possesses exactly four solutions. To address this discrepancy, we present the following proposition.
@@ 4. Case $B\neq 0$
Suppose that conditions $(C1)$ and $(C2)$ do not hold. Suppose that $h_{1}=0$ and $BC^{q}+B^{q}D\neq 0.$
Let $q$ be large enough. If $f_{A,B,C,D,E}(x)$ is APN then one of the following conditions possibly holds
@@ 5. $B\neq 0$ and (C3) does not hold
Solutions of $\overline{G}(X_{0},Z_{0},Z_{1})=0$ not contained in the forbidden hyperplanes can arise only from the factors $AD^{q}Z_{0}+AE^{q}Z_{0}^{2}+Z_{1}$ and $Z_{1}^{3}+AD^{q}Z_{0}Z_{1}^{2}+DZ_{0}^{2}Z_{1}+AZ_{0}^{3}$.
Let us consider the factor
@@ 6. Computational verification and discovery of new APN classes
To assess the diversity of discovered functions, we performed complete CCZ-equivalence classification using Magma implementations. For each field size, we tested pairwise CCZ-equivalence among all discovered APN functions to obtain exact counts of inequivalent classes.
@@
    For $q=2$ (exhaustive): $390$ APN functions in exactly $1$ CCZ-equivalence class, all CCZ-equivalent to the Budaghyan-Carlet (BC) family [3].
@@
    For $q=4$ (exhaustive): $28,170$ APN functions in exactly $1$ CCZ-equivalence class, all CCZ-equivalent to the BC family.
@@
    For $q=8$ ($120,000$ random candidates): $218$ APN functions in exactly $3$ distinct CCZ-equivalence classes, all outside of the BC family.
@@
    For $q=16$ ($200,000$ random candidates): $34$ APN functions in exactly $2$ distinct CCZ-equivalence classes (verified by complete pairwise testing), both outside of the BC family.
Representatives for each class appear in Tables LABEL:tab:ccz_F4LABEL:tab:ccz_F256. Remarkably, none of the discovered APN functions are permutations.
@@
The enumeration and classification were performed using SageMath implementations of the algorithms described above, followed by complete CCZ-equivalence testing using Magma. Complete computational code and output files are available at [11]. ∎
@@
Yet within this rarefied landscape, we find an interesting pattern. For the exhaustive searches at $q\in\{2,4\}$, all discovered APN functions belong to a single CCZ-equivalence class—the known Budaghyan-Carlet family. This suggests that for small fields, the BC construction captures the entire APN behavior within Dillon’s hexanomial framework. However, for larger fields $q\in\{8,16\}$, we discover new CCZ-equivalence classes outside the BC family, with $3$ classes for $q=8$ and $2$ classes for $q=16$. This validates that Dillon’s hexanomial family does contain inequivalent APN functions beyond the BC construction, though they appear only in larger fields.
The universal absence of permutations among all tested APN hexanomials is particularly striking. This distinguishes Dillon’s family from other APN constructions where permutation polynomials exist (albeit on odd dimensions), suggesting these hexanomials possess structural features fundamentally incompatible with bijectivity. Understanding this phenomenon could provide insight into the relationship between the APN property and injectivity in polynomial mappings over finite fields.